Which of the following sets Sj , (j = 1, . . . , 4) are NOT a neighbor-
hood of the given point a.
(1) S1 = (−2, 3), a = 1; (2) S2 = [−3, ∞), a = 2.
(3) S3 = {x: |x − 1| ≤ 2} , a = 3; (4) S4 = {x: |x − 3| ≥ 2} , a = 1.
1
Expert's answer
2021-09-30T02:16:34-0400

ANSWER: "S_3,S_4" are NOT a neighborhood of the given point a.

EXPLANATION.

Let "a\in S" and there exist "\varepsilon >0" such that "\left( a-\varepsilon ,a+\ \varepsilon \right) \subseteq S" . By the definition, "S" is a neighborhood of the point a.

Hence:

(1) "\left\{ 1 \right\} \in (-2,3)={ S }_{ 1 }" . Let "\u03b5=1," "\left( a-1,a+1 \right) =(1-1,1+1)=(0,2)\subset (-2,3)={ S }_{ 1 }\quad" ."S_1" is a neighborhood of the point a=1.

(2) "\left\{ 2 \right\} \in [-3,\infty)={ S }_{ 2 }" . Let "\u03b5=1," "\left( a-1,a+1 \right) =(2-1,2+1)=(1,3)\subset [-3,\infty )={ S }_{ 2 }\quad" ."S_2" is a neighborhood of the point a=2.

(3) "{ \left\{ 3 \right\} \in S }_{ 3 }=\left\{ x:|x-1|\le 2 \right\} =\left\{ x:-2\le x-1\ \le 2 \right\} =\left\{ x:-1\le x \ \le 3 \right\}" .

For all "\varepsilon >0\quad \left( a-\varepsilon ,a+\varepsilon \right) \ =(3-\varepsilon ,3+\varepsilon )\supset (3\ ,3+\varepsilon )\ \ , (3\ ,3+\varepsilon )\cap \ { S }_{ 3 }=\emptyset" . So, "S_3" is not a neighborhood of the point a=3.

(4) "{ \left\{ 1 \right\} \in S }_{ 4 }=\left\{ x:|x-3\ |\ge 2 \right\} =\left\{ x:\ \ x-3\le -2\ \ \right\} \cup \left\{ x:2\le x-3\ \ \right\} =\left\{ x:\ \ x\quad \le 1\ \ \right\} \cup \left\{ x:x\ge 5\ \ \ \right\} ." For all "\varepsilon >0\ \left( a-\varepsilon ,a+\varepsilon \right) \ =(1-\varepsilon ,1+\varepsilon )\supset (1\ ,1+\varepsilon )\quad" and "(1\ ,1+\varepsilon )\cap \left( 1,5 \right)" is not an empty set. Therefore, "(1\ ,1+\varepsilon )" is not completely contained in "S_4" . So, "S_4" is not a neighborhood of the point a=1.